OpenStudy (anonymous):

Jean says that you can take the square root of each side of an equation. Therefore, she decides that (x - 2)2 + (y + 3)2 = 36 and (x - 2) + (y + 3) = 6 are equuivalent equations. Marco says that the equations are not equivalent. Who is right? Explain your answer.

OpenStudy (anonymous):

\[\sqrt{a^2+b^2}\neq a+b\]

OpenStudy (anonymous):

\[\Large5=\sqrt{25}=\sqrt{16+9}\neq \sqrt{16}+\sqrt{9}\neq 4+3\neq 7\] because \[\LARGE 5\neq 7 \]

OpenStudy (anonymous):

so which one is right

OpenStudy (anonymous):

mine one is just an example of what satellite showed you !

OpenStudy (anonymous):

so how do i go about solving this equation